Each business day, on average, a company writes checks totaling $13,600 to pay its suppliers. The usual clearing time for the checks is four days. Meanwhile, the company is receiving payments from its customers each day, in the form of checks, totaling $24,600. The cash from the payments is available to the firm after two days.

a.

Calculate the company’s disbursement float, collection float, and net float. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

b. If the collected funds were available in one day instead of two, what would be the company's disbursement float, collection float, and net float? (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answers to the nearest whole number, e.g., 32.)

QEpok1LqJuu4iedJkVBUVRlMVCFHokoyeOOeaYY+
Answer #1

(a)

Each business day companies writing checks to suppliers of $13,600 and such checks takes 4 days time to clear .

Disbursement Float = Average amount of check* time to clear

= 13600*4

= 54400$

Collection Float =(-24600*2)

= -49200$

Net Float = Disbursement Float + collection Float

= 54400-49200

= 5200

(b)

Each business day companies writing checks to suppliers of $13,600 and such checks takes 4 days time to clear .

Disbursement Float = Average amount of check* time to clear

= 13600*4

= 54400$

Collection Float =(-24600*1)

= -24600$

Net Float = Disbursement Float + collection Float

= 54400-24600

= 29800$

I hope my efforts will be fruitful to you...?